Problem 31.37 A6.0 A current is charging a 1.4 -cm-diameter parallel plate capacitor. Part A What is the magnetic field strength at a point 2.2 mm radially from the center of the wire leading to the capacitor? Express your answer to two significant figures and include the appropriate units. Ba Value T Submit My Answers Give Up Part B What is the magnetic field strength at a point 2.2 mm radially from the center of the capacitor? Express your answer to two significant figures and include the appropriate units. B- Value T Subm My Answers Give Up Eeedback Continue
Answer #1

a) B = mu0 I / 2 pi r = (4 pi * 10-7 * 6) / (2 pi * 2.2 * 10-3)

= 5.45 * 10-4 T

b) B = mu0 / 2 pi * [r/R2] * dq/dt

= (4 pi * 10-7 * 2.2 * 10-3 * 6) / (2 pi * (0.7 * 10-2)2)

= 5.39 * 10-5 T